5. (a) Use the standard reduction potentials at 25° C in Table 18.1 in Tro, Fridgen and Shaw, and calculate the standard emf E° of an electrochemical cell described by the following reaction: Pb + Mn2+ --> Pb2+ + Mn (b) What is the value of E if the concentration of Mn2+ is 6.20 x 10-3 and the concentration of Pb2+ is 7.73 x 10-2? (c) If E = -0.89 V and Mn2+ = 0.10 M, what is the concentration of Pb2+?
Answer #1

Answer:-

This question is solved by using the simple concept of electrochemistry which involves oxidation and reduction using the Nernst equation.

The answer is given in the image,

Anwesend a Pb & Mn 2P - Pb²+ + M2 Pb Pb2+2e= E=+0.13 Mn?++2e-sMn = ° = -1.18 Eiel = 0,13 -1•18 =-1.05V bu, CMn 2+] = 6-20X103